OpenStudy (anonymous):

What two numbers, when multiplied equal -36 and when added equal -5.

OpenStudy (anonymous):

And it could be prime

OpenStudy (anonymous):

xy=-36, x+y=-5, y(-y-5)=-36, -y^2-5y+36=0, y^2+5y-36=0

OpenStudy (anonymous):

then use quadratic equation to find y and plug in to find x

OpenStudy (anonymous):

xy = -36 x+y = -5 x = -5 -y y(-5-y) = -36 -5y-y^2 = -36 y^2+5y-36 = 0 solve for Y and find X

OpenStudy (anonymous):

thnx